OpenStudy (anonymous):

Use a binomial series for f(x)= 1/(x+1)^4

OpenStudy (blockcolder):

Write that as \(\large f(x)=(1+x)^{-4}\) so that a series representation is: \[\large f(x)=\sum_{n=0}^{\infty} \binom{-4}{n} x^n\ \text{for } |x|\leq 1\]

OpenStudy (anonymous):

Thanks!

OpenStudy (blockcolder):

No problem. :D